a seller claimed that her lip tint has a mean organic content of 90%.A rival seller asked 60 users of that lip tint and found that it has a mean organic content of 85% with a standard deviation of 5%.Test the claim at 1% level of significance and assumes that the population is approximately normally distributed


1
Expert's answer
2021-06-01T16:10:10-0400

Hypothesized Population Mean "\mu=90"

Sample Standard Deviation "s=5"

Sample Size "n=60"

Sample Mean "\bar{x}=85"

Significance Level "\alpha=0.01"


Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu\geq90"

"H_1: \mu<90"

This corresponds to left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.01,"

"df=n-1=59" ​​degrees of fredom, and the critical value for two-tailed test i"t_c=2.661759."

The rejection region for this left-tailed test is "R=\{t:t<-2.391229\}."


The "t" - statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{85-90}{5/\sqrt{60}}\approx-7.745967"

Since it is observed that "t=-7.745967<-2.391229=t_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is lessthan "90," at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for left-tailed, the significance level "\alpha=0.01, df=59, t=-4.47214," is "p<0.00001," and since "p<0.00001<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than "90," at the "\alpha=0.01" significance level.